OpenStudy (anonymous):

Will give fan and medal! The area of a rectangular piece of land is 260 square meters. If the length of the land was 5 meters less and the width was 2 meters more, the shape of the land would be a square. Part A: Write an equation to find the width (x) of the land. Show the steps of your work. Part B: What is the width of the land in meters? Show the steps of your work.

OpenStudy (anonymous):

we have : lw = 260 -------1 and if the length of the land was 5 meters less and the width was 2 meters more, the shape of the land would be a square then : l-5 = w + 2 ==> l = w +7-------2 by plugging 2 in 1 ---> w^2 +7w = 260 -------> (w-13)(x+20)=0 w = - 20 neglected , then w = 13